The average monthly salary of 500 teachers of Candelaria East district is


P26,500. If the monthly salaries of the teachers are normally distributed with


standard deviation of P2,500, approximately how many teachers in the district


are earning more than P30,000 monthly?



1
Expert's answer
2022-04-01T04:44:26-0400

z="\frac{x-sample\space mean }{s/\sqrt n}"

n=500

sample mean =26500

standard deviation =2500

x=30000

P(more than 30000)=?

P(x>30000)=1-P(x"\le" 30000)

z="\frac{30000-26500}{2500/\sqrt{500}}"

z=31.30

p(z"\le"31.30)=1

P(x>30000)=0

Thus none of the teachers earn more than P 30000 monthly